Phosphorus reacts with oxygen to form diphosphorus pentoxide, P2O5 .



4P(s)+5O2(g)⟶2P2O5(s)



How many grams of P2O5 are formed when 2.17 g of phosphorus reacts with excess oxygen? Show the unit analysis used for the calculation by placing the correct components into the unit-factor slots.

1
Expert's answer
2022-03-19T02:43:36-0400

Moles of phosphorus are

Moles= 2.17/31= 0.07moles

Moles P2O5 are

(1/3)× 0.07 = 0.023moles

Mass= 284× 0.023

Mass= 6.532g